OpenStudy (anonymous):

Every time x increases by 1, y decreases by 3, and when x is -6, y is -3. Which function models the relationship between x and y? a) y= -6x - 21 b) y= -6x - 3 c) y= -3x - 3 d) y= -3x - 21

OpenStudy (anonymous):

It's b right ?

OpenStudy (anonymous):

Every time x increases by 1, y decreases by 3 --> this tells you something about the slope, because the slope is [change of y]/[change of x] when x is -6, y is -3 ---> this together with the slope allows you to figure out the starting position (the '+b' part. as in y=ax+b)

OpenStudy (anonymous):

so, based on that, what do you think the slope is, to start with?

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

i mean -6

OpenStudy (anonymous):

yeah -6

OpenStudy (anonymous):

Every time x increases by 1, y decreases by 3

OpenStudy (anonymous):

so \[slope = \frac{ \Delta y }{ \Delta x } = \frac{ -3 }{ 1 }\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

im still lost here does the -6 and -3 have to plugged in ?

OpenStudy (anonymous):

y=ax+b where a is the slope. you know the coordinates x=-6 and y=-3 -3 = -3 (-6) + b solve b and you'll know the answer

OpenStudy (anonymous):

So the answer is c?

OpenStudy (anonymous):

i think you missed a sign somewhere

OpenStudy (anonymous):

I think im doing it wrong

OpenStudy (anonymous):